# PowerInfer: Fast Large Language Model Serving with a Consumer-grade GPU

### v1

@@ Abstract
This paper introduces PowerInfer, a high-speed Large Language Model (LLM) inference engine on a personal computer (PC) equipped with a single consumer-grade GPU. The key underlying the design of PowerInfer is exploiting the high locality inherent in LLM inference, characterized by a power-law distribution in neuron activation. This distribution indicates that a small subset of neurons, termed hot neurons, are consistently activated across inputs, while the majority, cold neurons, vary based on specific inputs. PowerInfer exploits such an insight to design a GPU-CPU hybrid inference engine: hot-activated neurons are preloaded onto the GPU for fast access, while cold-activated neurons are computed on the CPU, thus significantly reducing GPU memory demands and CPU-GPU data transfers. PowerInfer further integrates adaptive predictors and neuron-aware sparse operators, optimizing the efficiency of neuron activation and computational sparsity. Evaluation shows that PowerInfer attains an average token generation rate of 13.20 tokens/s, with a peak of 29.08 tokens/s, across various LLMs (including OPT-175B) on a single NVIDIA RTX 4090 GPU, only 18% lower than that achieved by a top-tier server-grade A100 GPU. This significantly outperforms llama.cpp by up to 11.69$\times$ while retaining model accuracy.
1 Introduction
@@
Generative large language models (LLMs) have garnered attention for their remarkable capabilities in creative writing, advanced code generation, and sophisticated natural language processing tasks [5, 49, 42]. These models, widely deployed in data centers equipped with high-end and expensive server-grade GPUs, have significantly influenced our daily lives and work practices. Meanwhile, there is an emerging trend of running LLMs on more accessible local platforms, particularly personal computers (PCs) with consumer-grade GPUs. This evolution is driven by the need for enhanced data privacy [25], model customization [22], and reduced inference costs [42]. In contrast to data-center deployments, which prioritize high throughput [18, 37, 47], local deployments focus on low latency in processing small batches.
@@
Nonetheless, deploying LLMs on consumer-grade GPUs presents significant challenges due to their substantial memory requirements. LLMs, typically functioning as autoregressive Transformers, sequentially generate text token-by-token, each needing to access the entire model consisting of hundreds of billions of parameters. Therefore, the inference process is fundamentally constrained by the GPU’s memory capacity. This limitation is particularly acute in local deployments where the processing of individual requests (often just one at a time) [6] leaves minimal opportunity for parallel processing.
@@
Existing approaches to such memory issues include model compression and offloading. Compression techniques like quantization [12, 46], distillation [48], and pruning [23] reduce the model size. However, even deeply compressed models remain too large for consumer-grade GPUs. For instance, an OPT-66B model with 4-bit precision demands approximately 40GB of memory just to load its parameters [20], exceeding the capacity of even high-end GPUs like the NVIDIA RTX 4090. Model offloading, which partitions the model between GPU and CPU at the Transformer layer level [3, 37, 14]. State-of-the-art systems like llama.cpp [14] distribute layers between CPU and GPU memories, leveraging both for inference, thus reducing the GPU resources required. However, this method is hindered by the slow PCIe interconnect and the CPUs’ limited computational capabilities, resulting in high inference latency.
@@
In this paper, we argue that the key reason for memory issues in LLM inference is the locality mismatch between hardware architecture and the characteristics of LLM inference. Current hardware architectures are designed with a memory hierarchy optimized for data locality. Ideally, a small, frequently accessed working set should be stored in the GPU, which offers higher memory bandwidth but limited capacity. In contrast, larger, less frequently accessed data are better suited for CPUs, which provide more extensive memory capacity but lower bandwidth. Nevertheless, the vast volume of parameters required for each LLM inference iteration leads to a working set that is too large for a single GPU, thus impeding efficient locality exploitation.
@@
We have observed that LLM inference inherently exhibits high locality. Specifically, during each inference iteration, a limited number of neurons [^1] This paper defines a neuron as a specific row/column in a weight matrix. are activated, significantly influencing the outcome of token inference. These activations, which are input-specific, can be accurately predicted during runtime. For example, in the OPT model, less than 10% of the elements in the activation map are non-zero, and these can be predicted with more than 93% accuracy at runtime [21]. Notably, neuron activation in an LLM follows a skewed power-law distribution: a small subset of neurons consistently contributes to the majority of activations (over 80%) across various inputs (hot-activated), while the majority are involved in the remaining activations, which are determined based on the inputs at runtime (cold-activated).
@@
Building on the locality insights, we introduce PowerInfer, an efficient LLM inference system optimized for local deployments using a single consumer-grade GPU. The key idea of PowerInfer is to exploit the locality in LLM inference by assigning the minor hot neurons to the GPU, while cold neurons, which constitute the majority, are managed by the CPU. PowerInfer preselects and preloads hot-activated neurons onto the GPU offline and leverages online predictors during runtime to identify activated neurons. This approach allows the GPU and CPU to independently process their respective sets of neurons, thereby minimizing the need for costly PCIe data transfers.
@@
Second, leveraging LLM sparsity requires the use of sparse operators. Conventional libraries like cuSPARSE [30] are not optimal due to their general-purpose design, which includes tracking each non-zero element and converting dense matrices into sparse formats [45, 51]. In contrast, PowerInfer designs neuron-aware sparse operators that directly interact with individual neurons, thereby bypassing operations on entire matrices. This approach enables efficient matrix-vector multiplication at the neuron level and removes the need for specific sparse format conversions.
@@
The online inference engine of PowerInfer was implemented by extending llama.cpp with an additional 4,200 lines of C++ and CUDA code. Its offline component, comprising a profiler and a solver, builds upon the transformers framework [44] with approximately 400 lines of Python code. PowerInfer is compatible with various popular LLM families, including OPT (7B-175B), LLaMA (7B-70B), and Falcon-40B, and supports consumer-grade GPUs like the NVIDIA RTX 4090 and NVIDIA RTX 2080Ti.
Performance evaluation reveals that PowerInfer, when deployed on a PC equipped with a single NVIDIA RTX 4090 GPU, delivers an average generation speed of 13.20 tokens/s for quantized models and 8.32 tokens/s for non-quantized models, maintaining model accuracy. These results significantly surpass llama.cpp’s performance, exhibiting up to 8.00× and 11.69× improvements for quantized and non-quantized models, respectively. Significantly, the inference speed achieved on an NVIDIA RTX 4090 GPU (priced at approximately \$2,000) is only 18% slower compared to the performance on a top-tier A100 GPU (costing around \$20,000) that can fully accommodate the model. PowerInfer’s source code is publicly available at https://github.com/SJTU-IPADS/PowerInfer.
2 Background and Motivation
Figure 1: The inference procedure of an LLM.
2.1 LLM Inference & Architecture
LLM inference, an autoregressive model, generates each token based on previous ones. The process, illustrated in Figure 1, starts with a prompt (e.g., "I love reading") and unfolds in two phases: first, the prompt phase outputs an initial token ("OSDI"), then the generation phase sequentially produces tokens until a maximum limit or an end-of-sequence (<EOS>) token is reached. Each token generation, an inference iteration, requires running the full LLM model.
Figure 2: The architecture of a Transformer layer and how neurons are sparsely activated in FC1 and FC2 layers due to the ReLU function. The neurons that are activated are represented as green rows or columns encircled by red lines. The output vector from FC1 is then supplied to FC2 as its input vector.
@@ 2.1 LLM Inference & Architecture
In Figure 2 (right), the MLP block’s layers, FC1 and FC2, generate vectors through matrix multiplication. Each output element comes from the dot product of an input vector and a neuron (a row/column in a weight matrix). Activation functions like ReLU [1] act as gates to selectively retain or discard values in a vector, influencing neuron activations in FC1 and FC2. For example, ReLU in this figure filters out negative values, allowing only positively valued neurons in FC1 to influence the output. These neurons, which contribute to the output, are considered activated in this paper. Similarly, these values also affect which neurons in FC2 are activated and involved in the computation of its output vector.
Activation Sparsity. Recent studies have revealed that LLM inference shows a notable sparsity in neuron activation [19, 21, 50]. For example, we observe that approximately 80% of neurons in the OPT-30B model remain inactivated during the inference. This phenomenon of activation sparsity exists in both self-attention and MLP blocks. In self-attention blocks, nearly half of the attention heads (neurons) make minimal contributions, leading to their high sparsity. The sparsity observed within the MLP blocks is primarily attributed to the characteristics of the activation functions.
@@
Crucially, the activation sparsity is input-specific, meaning that the activation of specific neurons is directly influenced by the current input and cannot be predetermined before the model’s inference iteration begins. While it is not feasible to know which neurons will be activated before the entire model runs, it is possible to predict neuron activations a few layers in advance within the ongoing model iteration. DejaVu [21], for instance, utilizes MLP-based predictors during inference, achieving a remarkable accuracy rate of at least 93% in predicting neuron activation.
2.2 Offloading-based LLM Serving
@@
Figure 3: Typical existing offloading solutions. (a) shows a GPU-centric approach, while (b) is the CPU-GPU hybrid offloading approach.
@@
Figure 4: Performance comparison and analysis for serving OPT-30B on NVIDIA RTX 4090 GPU. The yellow blocks refer to FlexGen, the gray blocks refer to DejaVu (UM) and the blue blocks refer to llama.cpp. (a) The Y-axis indicates execution time for one iteration and the X-axis represents batch sizes for input. (b) The Y-axis indicates the proportion of execution time, and the X-axis indicates batch sizes for input.
Current model compression techniques are inadequate for fitting large language models (LLMs) within resource-limited consumer-grade GPUs. In contrast, the offloading technique, which leverages the CPU’s additional computational and memory resources, presents a more viable solution for accommodating LLMs on such hardware. Figure 3 illustrates two main offloading approaches:
@@
GPU-centric offloading utilizes CPU memory to store portions of the model parameters that exceed the GPU’s capacity. During each iteration, as depicted in Figure 3a), it processes the parameters located in the GPU memory, transferring more from the CPU as needed. This strategy enables the inference of LLMs of varying sizes, provided that sufficient combined CPU memory and hard disk storage are available. FlexGen [37] is a typical example that adopts a zig-zag scheduling approach to prioritize throughput over latency, processing batches sequentially for each layer. Nonetheless, this method leads to substantial per-token latency in latency-sensitive scenarios (Figure 4a), mainly due to frequent data transfers between GPU and CPU, especially with batch sizes of one. Over 99.5% of processing time is consumed by transferring LLM weights from CPU to GPU, significantly impacting overall latency, as illustrated in Figure 4b.
@@
DejaVu [21] accelerates LLM inference by using activation sparsity. It selectively processes only those neurons that are predicted to be activated (called predicted neurons for brevity), while bypassing the inactivated ones. However, this approach, initially designed for data center inference, struggles on consumer-grade GPUs that cannot accommodate full-scale LLMs. The key challenge with DejaVu in such contexts stems from the need to frequently transfer activated neurons from the CPU to the GPU during runtime. For LLMs like OPT-30B that exceed GPU memory limits, DejaVu [^2] Since DejaVu only works for GPU, we modified it by using NVIDIA Unified Memory (UM) [29] to fetch parameters from CPU memory., albeit reducing the computational load on the GPU, is constrained by the data transfer procedure (Figure 4a). Consequently, as shown in Figure 4a, DejaVu experiences significant inference latency, comparable to that of FlexGen.
@@
Hybrid offloading distributes model parameters between GPU and CPU, splitting them at the Transformer layer level as shown in llama.cpp [14] (Figure 3b). The CPU processes its layers first, then sends intermediate results to the GPU for token generation. This offloading method reduces inference latency to around 600ms (Figure 4a) by minimizing data transfer and mitigating slow PCIe bandwidth.
@@
However, hybrid offloading still faces the locality mismatch issue, leading to suboptimal latency. Each inference iteration accesses the entire model, resulting in poor locality for hierarchical GPU-CPU memory structures. GPUs, while computationally powerful, are constrained by memory capacity. For instance, a 30B-parameter model on a 24GB NVIDIA RTX 4090 GPU means only 37% of the model is on the GPU, shifting most computational tasks to the CPU. The CPU, with higher memory but lower computational power, ends up handling 98% of the total computational load (Figure 4b).
3 Insights into Locality in LLM Inference
This section introduces our insights into locality in the LLM inference procedure, highlighting two distinctive features.
3.1 Insight-1: Power-law Activation

### v2

@@ Abstract.
This paper introduces PowerInfer, a high-speed Large Language Model (LLM) inference engine on a personal computer (PC) equipped with a single consumer-grade GPU. The key principle underlying the design of PowerInfer is exploiting the high locality inherent in LLM inference, characterized by a power-law distribution in neuron activation. This distribution indicates that a small subset of neurons, termed hot neurons, are consistently activated across inputs, while the majority, cold neurons, vary based on specific inputs. PowerInfer exploits such an insight to design a GPU-CPU hybrid inference engine: hot-activated neurons are preloaded onto the GPU for fast access, while cold-activated neurons are computed on the CPU, thus significantly reducing GPU memory demands and CPU-GPU data transfers. PowerInfer further integrates adaptive predictors and neuron-aware sparse operators, optimizing the efficiency of neuron activation and computational sparsity. The evaluation shows that PowerInfer significantly outperforms llama.cpp by up to 11.69$\times$ while retaining model accuracy across various LLMs (including OPT-175B) on a single NVIDIA RTX 4090 GPU. For the OPT-30B model, PowerInfer achieves performance comparable to that of a high-end server-grade A100 GPU, reaching 82% of its token generation rate on a single consumer-grade RTX 4090 GPU.
LLM serving, Sparsity
journalyear: 2024copyright: acmlicensedconference: ACM SIGOPS 30th Symposium on Operating Systems Principles; November 4–6, 2024; Texas, USAbooktitle: ACM SIGOPS 30th Symposium on Operating Systems Principles (SOSP ’24), November 4–6, 2024, Texas, USAjournalyear: 2024copyright: acmlicensedconference: ACM SIGOPS 30th Symposium on Operating Systems Principles; November 4–6, 2024; Austin, TX, USAbooktitle: ACM SIGOPS 30th Symposium on Operating Systems Principles (SOSP ’24), November 4–6, 2024, Austin, TX, USAdoi: 10.1145/3694715.3695964isbn: 979-8-4007-1251-7/24/11ccs: Computing methodologies Machine learning
1. Introduction
@@
Generative large language models (LLMs) have garnered significant attention for their remarkable capabilities in sophisticated natural language processing tasks (Brown et al., 2020; Zhang et al., 2022b; Touvron et al., 2023), especially in areas such as creative writing and advanced code generation. These models, widely deployed in data centers equipped with high-end and expensive server-grade GPUs, have significantly influenced our daily lives and work practices. Meanwhile, there is an emerging trend of running LLMs on more accessible local platforms (Xue et al., 2024; Chun et al., 2011), particularly personal computers (PCs) with consumer-grade GPUs. This evolution is driven by the need for enhanced data privacy (Martínez Toro et al., 2023), model customization (Lyu et al., 2023), and reduced inference costs (Touvron et al., 2023). In contrast to data-center deployments, which prioritize high throughput (Kwon et al., 2023; Sheng et al., 2023b; Yu et al., 2022), local deployments focus on low latency in processing small batches.
@@
Nonetheless, deploying LLMs on consumer-grade GPUs presents significant challenges due to their substantial memory requirements. LLMs, functioning as autoregressive Transformers, sequentially generate text token-by-token, each needing to access the entire model containing hundreds of billions of parameters. Therefore, the inference process is fundamentally constrained by the GPU’s memory capacity. This limitation is particularly acute in local deployments where the processing of individual requests (often one at a time) (Cai et al., 2023) leaves minimal opportunity for parallel processing.
@@
Compression techniques like quantization (Frantar et al., 2022; Xiao et al., 2023) and pruning (Ma et al., 2023a) can reduce the model size. However, even deeply compressed models remain too large for consumer-grade GPUs. For instance, an OPT-66B model with 4-bit precision demands approximately 40GB of memory just to load its parameters (Liu et al., 2023a), exceeding the capacity of even high-end GPUs like the NVIDIA RTX 4090.
Model offloading is another approach that partitions the model between GPU and CPU at the Transformer layer level (Aminabadi et al., 2022; Sheng et al., 2023b; Gerganov, 2023). State-of-the-art systems like llama.cpp (Gerganov, 2023) distribute layers between CPU and GPU memories, leveraging both for inference, thus reducing the GPU resources required. However, this method is hindered by the slow PCIe interconnect and the CPUs’ limited computational capabilities, resulting in high inference latency.
@@
In this paper, we argue that the key reason for memory issues in LLM inference is the locality mismatch between hardware architecture and the characteristics of LLM inference. Current hardware architectures are designed with a memory hierarchy optimized for data locality. Ideally, a small, frequently accessed working set should be stored in the GPU, which offers higher memory bandwidth but limited capacity. In contrast, larger, less frequently accessed data are better suited for CPUs, which provide more extensive memory capacity but lower bandwidth. Nevertheless, each LLM inference iteration requires accessing the entire set of model parameters whose total size is too large for a single GPU, thus showing no locality at all and thus impeding efficient locality exploitation.
Recent works have identified activation sparsity in LLM inference (Liu et al., 2023b; Mirzadeh et al., 2023; Zhang et al., 2024b). During each inference iteration, only a limited number of neurons [^1] This paper defines a neuron as a specific row/column in a weight matrix. are activated, significantly influencing token outputs. These sparse activations, which can be accurately predicted at runtime, allow for accelerated inference by computing only the activated neurons. However, the set of activated neurons varies across inputs and can only be determined at runtime, necessitating the entire model to be loaded into GPU memory. This requirement limits the approach’s applicability in local deployment scenarios with constrained GPU VRAM.
@@
Fortunately, we have observed that neuron activation in an LLM follows a skewed power-law distribution across numerous inference processes: a small subset of neurons consistently contribute to the majority of activations (over 80%) across various inputs (hot-activated), while the majority are involved in the remaining activations, which are determined based on the inputs at runtime (cold-activated). This observation suggests an inherent locality in LLMs with high activation sparsity, which could be leveraged to address the aforementioned locality mismatch.
@@
Building on this locality insight, we introduce PowerInfer, an efficient LLM inference system optimized for local deployments using a single consumer-grade GPU. The key idea of PowerInfer is to exploit the locality in LLM inference by assigning the minor hot neurons to the GPU, while cold neurons, which constitute the majority, are managed by the CPU. Specifically, PowerInfer exploits the locality in LLM inference through a two-step process: (1) PowerInfer preselects hot and cold neurons based on their statistical activation frequency, preloading them onto the GPU and CPU, respectively, during an offline phase. (2) At runtime, it employs online predictors to identify which neurons (both hot and cold) are likely to be activated for each specific input. This approach allows the GPU and CPU to independently process their respective sets of activated neurons, thereby minimizing the need for costly PCIe data transfers.
@@
Second, leveraging LLM sparsity requires the use of sparse operators. Conventional libraries like cuSPARSE (NVIDIA, 2023) are not optimal due to their general-purpose design, which includes tracking each non-zero element and converting dense matrices into sparse formats (Xia et al., 2023; Zheng et al., 2023b). In contrast, PowerInfer designs neuron-aware sparse operators that directly interact with individual neurons, thereby bypassing operations on entire matrices. This approach enables efficient matrix-vector multiplication at the neuron level and removes the need for specific sparse format conversions.
@@
The online inference engine of PowerInfer was implemented by extending llama.cpp with an additional 4,200 lines of C++ and CUDA code. Its offline component, comprising a profiler and a solver, builds upon the transformers framework (Wolf et al., 2020) with approximately 400 lines of Python code. PowerInfer is compatible with various popular LLM families, including OPT (7B-175B), LLaMA2 (7B-70B), and Falcon-40B, and supports consumer-grade GPUs like the NVIDIA RTX 4090 and NVIDIA RTX 2080Ti.
Performance evaluation reveals that PowerInfer, when deployed on a PC equipped with a single NVIDIA RTX 4090 GPU, delivers an average generation speed of 13.20 tokens/s for quantized models and 8.32 tokens/s for non-quantized models, maintaining model accuracy. These results significantly surpass llama.cpp’s performance, exhibiting up to 8.00× and 11.69× improvements for quantized and non-quantized models, respectively. Significantly, the inference speed achieved on an NVIDIA RTX 4090 GPU (priced at approximately \$2,000) is only 18% slower compared to the performance on a top-tier A100 GPU (costing around \$20,000) that can fully accommodate the model. PowerInfer’s code has been open sourced completely.
2. Background and Motivation
2.1. LLM Inference & Architecture
LLM inference, an autoregressive model, generates each token based on previous ones. The process starts with a prompt and unfolds in two phases: first, the prompt phase outputs an initial token, then the generation phase sequentially produces tokens until a maximum limit or an end-of-sequence (¡EOS¿) token is reached. Each token generation, an inference iteration, requires running the full LLM model.
Figure 1. The architecture of a Transformer layer and how neurons are sparsely activated in FC1 and FC2 layers due to the ReLU function. The neurons that are activated are represented as green rows or columns encircled by red lines. The output vector from FC1 is then supplied to FC2 as its input vector.
@@ 2.1. LLM Inference & Architecture
In Figure 1 (right), the MLP block’s layers, FC1 and FC2, generate vectors through matrix multiplication. Each output element comes from the dot product of an input vector and a neuron (a row/column in a weight matrix). Activation functions (like ReLU (Agarap, 2018) and SiLU (Ramachandran et al., 2017)) act as gates to selectively retain or discard values in a vector, influencing neuron activations in FC1 and FC2. For example, ReLU in this figure filters out negative values, allowing only positively valued neurons in FC1 to influence the output. These neurons, which contribute to the output, are considered activated in this paper. Similarly, these values also affect which neurons in FC2 are activated and involved in the computation of its output.
Activation Sparsity. LLM inference exhibits notable sparsity in neuron activation, a phenomenon observed in both self-attention and MLP blocks (Li et al., 2022; Liu et al., 2023b; Zhang et al., 2022a). In self-attention, nearly half of the attention heads contribute minimally, while in MLP blocks, sparsity is largely due to activation function characteristics. This MLP sparsity is particularly pronounced in ReLU-based architectures, with DejaVu (Liu et al., 2023b) reporting that approximately 80% of neurons in the OPT-30B model remain inactive during inference. Moreover, this phenomenon is not limited to ReLU-based models; it is also observed in architectures using other activation functions such as SwiGLU. For instance, Table 1 shows that LLaMA2-13B and Yi-34B, which employ SwiGLU, exhibit sparsity levels of 43% and 53% respectively. These findings align with prior research results from studies like CATS (Lee et al., 2024b) and ReLU2 (Zhang et al., 2024b).
Table 1. Average activation sparsity for various LLMs in the MLP blocks. For ReLU-based models, sparsity is the proportion of neurons with zero activation. For SwiGLU-based models, it’s the proportion of neurons that can be dynamically pruned with less than 1% impact on perplexity.
LLM Activation Function Sparsity
OPT-30B ReLU 97%
LLaMA2-13B SwiGLU 43%
Yi-34B SwiGLU 53%
@@
Moreover, it is possible to predict neuron activations a few layers in advance within the ongoing model iteration. Based on this observation, DejaVu (Liu et al., 2023b) employs MLP-based online predictors during inference and only processes the activated neurons, achieving over a 6x speedup while maintaining an impressive accuracy rate of at least 93% in predicting neuron activation. However, the activation sparsity is input-specific for each inference iteration, meaning that the activation of specific neurons is directly influenced by the current input and cannot be predetermined before the model’s inference iteration begins.
2.2. Offloading-based LLM Serving
@@
Figure 2. Typical existing offloading solutions. (a) shows a GPU-centric approach, while (b) is the CPU-GPU hybrid offloading approach.
@@
Figure 3. Performance comparison and analysis for serving OPT-30B on NVIDIA RTX 4090 GPU. The yellow blocks refer to FlexGen, the gray blocks refer to DejaVu (UM) and the blue blocks refer to llama.cpp. (a) The Y-axis indicates execution time for one iteration and the X-axis represents batch sizes for input. (b) The Y-axis indicates the proportion of execution time, and the X-axis indicates batch sizes for input.
The offloading technique, which leverages the CPU’s additional computational and memory resources, presents a more viable solution for accommodating LLMs that exceed the GPU’s memory capacity. In this section, we delve into the analysis of offloading systems to uncover the factors contributing to their sluggish performance. Figure 2 illustrates two main offloading approaches:
@@
GPU-centric offloading utilizes CPU memory to store portions of the model parameters that exceed the GPU’s capacity. During each iteration, as depicted in Figure 2a), it processes the parameters located in the GPU memory, transferring more from the CPU as needed. This strategy enables the inference of LLMs of varying sizes, provided that sufficient combined CPU memory and hard disk storage are available. FlexGen (Sheng et al., 2023b) is a typical example that adopts a zig-zag scheduling approach to prioritize throughput over latency, processing batches sequentially for each layer. Nonetheless, this method leads to substantial per-token latency in latency-sensitive scenarios (Figure 3(a)), mainly due to frequent data transfers between GPU and CPU. Over 99.5% of processing time is consumed by transferring LLM weights from CPU to GPU, significantly impacting overall latency, as illustrated in Figure 3(b).
@@
Although DejaVu (Liu et al., 2023b) leverages activation sparsity to accelerate LLM inference, this approach, originally designed for data center environments, faces challenges when applied to consumer-grade GPUs incapable of hosting full-scale LLMs. The key challenge with DejaVu in such contexts stems from the need to frequently transfer activated neurons from the CPU to the GPU during runtime. For LLMs like OPT-30B that exceed GPU memory limits, DejaVu [^2] Since DejaVu only works for GPU, we modified it by using NVIDIA Unified Memory (UM) (NVIDIA, 2021) to fetch parameters from CPU memory., albeit reducing the computational load on the GPU, is constrained by the data transfer procedure (Figure 3(a)). Consequently, as shown in Figure 3(a), DejaVu experiences significant inference latency, comparable to that of FlexGen.
@@
Hybrid offloading distributes model parameters between GPU and CPU, splitting them at the Transformer layer level (Figure 2b), with llama.cpp (Gerganov, 2023) as an example. The CPU processes its layers first, then sends intermediate results to the GPU for token generation. This offloading method reduces inference latency to around 600ms (Figure 3(a)) by minimizing data transfer and mitigating slow PCIe bandwidth. However, compared to the 45ms latency of the 30B model on A100, the speed is still too slow.
@@
Hybrid offloading still faces the locality mismatch issue, leading to suboptimal latency. Each inference iteration accesses the entire model, resulting in poor locality for hierarchical GPU-CPU memory structures. GPUs, while computationally powerful, are constrained by memory capacity. For instance, a 30B-parameter model on a 24GB NVIDIA RTX 4090 GPU means only 37% of the model is on the GPU, shifting most computational tasks to the CPU. The CPU, with higher memory but lower computational power, ends up handling 98% of the total computational time(Figure 3(b)).
3. Insights into Locality in LLM Inference
3.1. Insight-1: Power-law Activation